001 10.0 points To recharge a 12 V battery, a battery charger must move 3.8 x 105 C of charge from the negative terminal to the positive terminal How much work is done by the battery charger? Answer in units of J
Answer #1

The equation for work done to move charge from negative terminal to positive terminal is

W = q V

= 3.8* 10^5 ( 12 V)

=45.6* 10^5 J

or

= 4.56 * 10^6 J